∆ABC is right angled at B. BD is an altitude. If DC = 5 cm and AD = 40 cm, what is the value of BC (in cm)?

Grade:10

1 Answers

sahil
6 years ago
see in triangle abc with altitude d on ac solution:We get 2 other right angle triangle BDC and BDA both right angled at D.Now by pytha we have BC^2=5^2+DB^2 and in other AB^2=40^2+DB^2 and ac is 45 so in abc 45^2=AB^2+BC^2 =40^2+5^2+2DB^2we get DB^2=200now as early BC^2=5^2+DB^2=25+200=225=15^2SO BC =15CMif this was helpful please give approval